OpenStudy (lena772):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (anonymous):

y-y1=m(x-1) using the two points will give you the equation for the line, and since you already have the slope for the line, you should easily be able to solve.

OpenStudy (anonymous):

passing through the point means the points satisfies its equation

OpenStudy (lena772):

y+3=-6x+6

OpenStudy (anonymous):